Y=0.12x+0.76
0.11x-Y=0.85
1)Are these lines intersecting parellel or perpendicular?
2) Use the equation of the lines to determine if they are parellel.

Answer by mananth(16946)

Y=0.12x+0.76
0.11x-Y=0.85
re write the equation in slope intercept form
y=0.11x-0.85
...
compare the two lines.
if the lines are parallel the slopes should be same.
They are not so they are not parallel.
..
If the lines are perpendicular then the slopes should be negative reciprocal of each other. they are not.
so they are neither perpendicular.
..
solve the two equations
..
Y-0.12x=0.76
0.11x-Y=0.85
...
add the two equations
-0.01X=1.61
X=-161
..
plug value of x to get y
Y-0.12*-161=0.76
Y=-18.56
...
m.ananth@hotmail.ca